Infinite sheet carrying surface charge density sig
Answer #1

a)

Field due to infinite sheet carrying surface,

E = sigma / 2e0

E = (1 x 10^-6) / (2 x 8.854 x 10^-12 ) =56471.65 N/C towards the sheet


and force = q E = 8 x pi x 10^-6 x 56471.65

= 1.42 N toward the sheet


b) field due to +ve point charge is always away from it.

so point where field is zero will be x > 1

suppose coordinate is x .

then kq/d^2 - 56471.65 = 0

(9 x 10^9 x 8x pi x 10^-6) / ((x - 1)^2) = 56471.65

x - 1 = 2 m

x = 3 m ............Ans